# ConVerSum: A Contrastive Learning based Approach for Data-Scarce Solution of Cross-Lingual Summarization Beyond Direct Equivalents

### v1

@@ 4.1 Training Setup
4.2 Training Consideration
We have two things in our consideration for training ConVerSum - (i) training utilizing a large-scale monolingual dataset of high-resource languages and (ii) training utilizing a significantly smaller monolingual dataset of low-resource languages to demonstrate the power of contrastive learning in ConVerSum for both large-scale and small-scale datasets. The ConVerSum’s performance is analyzed on a cross-lingual dataset to verify its robustness.
4.3 Datasets
According to the consideration mentioned in Section 4.2. We choose CNN/DailyMail [32, 33] for the high-resource large-scale monolingual dataset and XLSUM [34] for low-resource significantly smaller monolingual datasets. We choose the monolingual Bengali, Thai, Burmese, and Tigrinya language datasets for our task from XLSUM. Finally, we perform a performance analysis of our model on CrossSum [6], a large-scale CLS dataset comprising 1.68 million article-summary samples in 1,500+ language pairs. The dataset for training and verification of the model is shown in Table 1. The dataset statistics for monolingual datasets for training are shown in Table 2.
Table 1: Datasets for Training and Verification
@@ 4.3 Datasets
Table 2: Training, Validation, and Test Splits for Monolingual Datasets for Training(Rsrc=Resource, Ds=Dataset)
Rsrc-Ds Monolingual Data Train Val Test
@@
4.4 Baseline Model
We use the mT5 [25] and FLAN-T5 [31] models as baselines for our tasks. These models are selected for their effectively handled multilingual data and demonstrated state- of-the-art performance in tasks related to NLP, such as text summarization.
4.5 Evaluation Metrics
We use LaSE [6] and BERTScore [35]. LaSE is specifically designed for evaluating CLS considering meaning similarity, language confidence and length penalty. BERTScore exploits the pre-trained contextual embeddings from BERT [36] to identify words in candidate and reference sentences based on cosine similarity.
4.6 ConVerSum: Training on High-Resource Large Scale Dataset
@@
To generate candidate summaries in different languages, we use mT5 checkpoint [^5] https://huggingface.co/csebuetnlp/mT5_m2m_crossSum_enhanced available in Huggingface [^6] https://huggingface.co/. For diversity sampling using diverse Beam search, we employ 8 groups, yielding 8 candidate summaries. The maximum length of the generated candidate summaries is specified at 80 words. Using an identical length guarantees uniformity between summaries in different languages. This uniformity makes the evaluation process easier and allows for fair comparisons. Moreover, languages differ in terms of word length and sentence structure. The summary word restriction guarantees all summaries are reasonably equal in terms of their content and depth, independent of language. For validation and test sets, we also generate the candidate summaries. To encode reference summaries, candidate summaries, and documents, we initiate the evaluation function with the XLM-RoBERTa checkpoint [^7] FacebookAI/xlm-roberta-base as it learns the inner representation of 100 languages. Then, we apply the contrastive learning objective described in Section 3. For training, Adam [37] optimizer is used with learning rate scheduling. The batch size is set to 2 for candidate generation and 4 for model training after trying 32, 16, 8 and 4 because of our GPU limitations. This indicates that we always choose the maximum batch size allowed by our GPU memory. We evaluate the performance of ConVerSum on the validation set at every 1000 steps using the LaSE. We set 15 epochs for training.
@@
4.6.1 Performance Comparison of ConVerSum with Baselines on CrossSum Dataset
@@
Table 3: Performance comparison of (English source and other target) baseline and ConVerSum based on mT5 on CrossSum
@@
Table 4: Performance comparison of (English source and Others target) baseline and ConVerSum based on Flan-T5 on CrossSum
@@
Table 5: Performance comparison of (Others source and English target) baseline and ConVerSum based on mT5 on CrossSum
@@
Table 6: Performance comparison of (Others source and English target) baseline and ConVerSum based on Flan-T5 on CrossSum
@@
Table 7: Performance comparison of (Others source and Others target) baseline and ConVerSum based on mT5 on CrossSum
@@
Table 8: Performance comparison of (Others source and Others target) baseline and ConVerSum based on Flan-T5 on CrossSum
@@
4.6.2 Performance Comparison of ConVerSum and LLM for Low Resource Languages
@@
Evaluating the performance of the state-of-art LLMs on various datasets and numerous NLP tasks is a trending research [38, 39]. Inspired by this trend, in this section, we compare the performance of ConVerSum to cutting-edge LLMs - ChatGPT-3.5, ChatGPT-4o, and Gemini 1.5 Pro to measure our ConVerSum’s efficiency in generating high-quality summaries in a variety of low-resource languages. To assess LLMs’ proficiency, we prompt them to rate their level of confidence between 1 to 10 in generating accurate cross-lingual summaries in 45 different languages. Responses are repeatedly generated to ensure clarity. All LLMs rate their capability as low (1-4) for CLS on low-resource languages. We choose three language pairs mentioned in Table 9 based on their ratings for this experiment.
{mdframed}
@@
[linecolor=black, linewidth=2pt, leftmargin=1em, rightmargin=1em, innertopmargin=1em, innerbottommargin=1em] "How confident are you to generate high-quality cross-lingual summary concisely and informatively for low-resource languages? Find the list of languages - Amharic, Arabic, Azerbaijani, Bengali, Burmese, Chinese_simplified, Chinese_traditional, English, French, Gujarati, Hausa, Hindi, Igbo, Indonesian, Japanese, Kirundi, Korean, Kyrgyz, Marathi, Nepali, Oromo, Pashto, Persian, Pidgin, Portuguese, Punjabi, Russian, Scottish_gaelic, Serbian_cyrillic, Serbian_latin, Sinhala, Somali, Spanish, Swahili, Tamil, Telugu, Thai, Tigrinya, Turkish, Ukrainian, Urdu, Uzbek, Vietnamese, Welsh, Yoruba. Rate your confidence level for cross-lingual summarization on a scale of 1 to 10 for the given languages."
@@
Then, we use manual prompting for LLMs which is free of cost. For our selected language pairs, we provide the following prompt where the source document is provided from CrossSum [6] test set and extract the responses manually from LLMs to measure the LaSE and BERTScore and compare those with ConVerSum which are mentioned in Table 9. {mdframed}[linecolor=black, linewidth=2pt, leftmargin=1em, rightmargin=1em, innertopmargin=1em, innerbottommargin=1em] "Summarize the given text in target language, preferably in 80 words, concisely and informative. .. (source document is given here)"
Table 9: Some Performance Measures on LLM vs Our Proposed Model for Low-resource Languages
@@
4.7 ConVerSum: Training on Low-Resource Language Datasets
This section emphasizes how we develop ConVerSum for CLS on XLSUM [34] Dataset as we mentioned in Section 4.3 to observe how ConVerSum performs on the low-resource small-scale datasets. We choose Bengali, Thai, Burmese and Tigrinya from the XLSUM dataset. As the results using mT5 outperform Flan-T5 in Section 4.6.1, we choose only using mT5 in the case discussed in this section. Therefore, all training setups are kept the same setup mentioned in Section 4.6.
4.7.1 Performance Comparison of ConVerSum with Baselines on CrossSum Dataset
Table 10: Performance Comparison of ConVerSum Trained on XLSUM-Bengali with Baseline on CrossSum Dataset
Baseline (mT5) ConVerSum
Source-Target LaSE BERTScore LaSE BERTScore
Bengali-English 0.5168 0.8263 0.5148 0.8257
Bengali-Arabic 0.5650 0.8229 0.5535 0.8205
Bengali-Marathi 0.6197 0.8493 0.6357 0.8545
Ukrainian-Bengali 0.6110 0.8493 0.6114 0.8470
Hausa-Bengali 0.5831 0.8493 0.6846 0.8629
Table 11: Performance Comparison of ConVerSum Trained on XLSUM-Thai with Baseline on CrossSum Dataset
Source-Target Baseline(mT5) ConVerSum
LaSE BERTScore LaSE BERTScore
Thai-English 0.5078 0.8298 0.5223 0.8315
Thai-Bengali 0.5937 0.8445 0.6447 0.8487
Thai-Marathi 0.5998 0.8411 0.6677 0.8465
Uzbek-Thai 0.6672 0.9305 0.6672 0.9390
Amharic-Thai 0.3783 0.8340 0.4310 0.8268
Table 12: Performance Comparison of ConVerSum Trained on XLSUM-Burmese with Baseline on CrossSum Dataset
Source-Target Baseline(mT5) ConVerSum
LaSE BERTScore LaSE BERTScore
Burmese-English 0.5394 0.8287 0.5422 0.8283
Burmese-Bengali 0.6146 0.8400 0.5973 0.8448
Burmese-Marathi 0.4247 0.8371 0.6821 0.8648
Uzbek-Burmese 0.4918 0.8335 0.4794 0.8431
Amharic-Burmese 0.6345 0.8689 0.6054 0.8599
Table 13: Performance Comparison of ConVerSum Trained on XLSUM-Tigrinya with Baselines on CrossSum Dataset
Baseline(mT5) ConVerSum
Source-Target LaSE BERTScore LaSE BERTScore
Tigrinya-English 0.4478 0.7935 0.3758 0.7937
Tigrinya-Arabic 0.4137 0.7837 0.5241 0.7966
Tigrinya-Marathi 0.4185 0.7762 0.6423 0.7916
Chinese_simplified-Tigrinya 0.4390 0.8288 0.5497 0.8288
Somali-Tigrinya 0.2241 0.7829 0.5537 0.7801
5 Result Analysis
5.1 For Training on High-Resource Large-Scale Dataset
From Table 3 we can say that ConVerSum performs better for LaSE for all source-target language pairs we choose which proves the effectiveness of our model in CLS. While calculating BERTScore, three cases - English to Chinese_simplified, Thai and Tigrinya, ConVerSum achieves almost similar performance as the baseline mT5. Table 5 shows that ConVerSum achieves better performance for LaSE for all of the source-target language pairs we experimented with except Bengali to English and Tigrinya to English though these two cases show almost similar performance to the baseline. While calculating BERTScore, three cases - Hindi-English, Bengali-English and Marathi-English, ConVerSum achieves almost identical performance as the baseline mT5. Table 7 illustrates that ConVerSum achieves better performance for LaSE for almost all of the source-target language pairs we experimented with. While calculating BERTScore, for three cases, ConVerSum achieves almost similar performance as the baseline mT5.
From Table 4, Table 6 and Table 8, it is clear that the performance of ConVerSum based on Flan-T5 and baseline Flan-T5 itself performs significantly lower than the ConVerSum of mT5 and also mT5 as a baseline itself for all combinations mentioned in Section 4.6.1.
5.2 For Training on Low-Resource Small-Scale Dataset
From Table 10, Table 11, Table 12 and Table 13 we can wrap up that ConVerSum outperforms baseline for most of the language pairs and only for a few instances, it shows almost identical performance to baseline.
5.3 ConVerSum vs LLMs

### v2

@@ 4.1. Training Setup
4.2. Training Consideration
We have two things in our consideration for training ConVerSum. The first is to train ConVerSum utilizing a large-scale monolingual dataset of high-resource languages. The second is to train ConVerSum utilizing a significantly smaller monolingual dataset of low-resource languages. The reason behind considering these two options is to demonstrate how ConVerSum works while training the model on a large-scale dataset of high-resource language and a significantly smaller dataset of low-resource language which proves the power of contrastive learning for both large-scale and small-scale datasets. For both cases, we analyze the performance of ConVerSum on a cross-lingual dataset to verify the model’s robustness.
4.3. Datasets
According to the consideration mentioned in Section 4.2. We choose CNN/DailyMail (Hermann et al., 2015; Nallapati et al., 2016) for the high-resource large-scale monolingual dataset and XLSUM (Hasan et al., 2021) for low-resource significantly smaller monolingual datasets. We choose the monolingual Bengali, Thai, Burmese, and Tigrinya language datasets for our task from XLSUM. Finally, we perform a performance analysis of our model on CrossSum (Bhattacharjee et al., 2023), a large-scale CLS dataset comprising 1.68 million article-summary samples in 1,500+ language pairs. The dataset for training and verification of the model is shown in Table 1. The dataset statistics for monolingual datasets for training are shown in Table 2. The dataset splits are publicly available and provided by the author and intended for efficient training, validation, and unbiased testing to ensure robust model performance across various languages and summarization tasks.
Table 1. Datasets for Training and Verification
@@ 4.3. Datasets
Table 2. Training, Validation, and Test Splits for Monolingual Datasets for Training
Resource-Dataset Monolingual Data Train Val Test
@@
4.4. Baseline Model
We use the mT5 (Xue et al., 2021) and FLAN-T5 (Chung et al., 2022) models as baselines for our tasks. These models are selected for their effectively handled multilingual data and demonstrated state- of-the-art performance in tasks related to NLP, such as text summarization.
4.5. Evaluation Metrics
We use LaSE (Bhattacharjee et al., 2023) and BERTScore (Zhang* et al., 2020). LaSE is specifically designed for evaluating CLS considering meaning similarity, language confidence and length penalty. BERTScore exploits the pre-trained contextual embeddings from BERT (Devlin et al., 2019) to identify words in candidate and reference sentences based on cosine similarity.
4.6. First Consideration: Training ConVerSum on High-Resource Large Scale Dataset
@@
To generate candidate summaries in different languages, we use mT5 checkpoint [^6] https://huggingface.co/csebuetnlp/mT5_m2m_crossSum_enhanced available in Huggingface [^7] https://huggingface.co/. For diversity sampling using diverse Beam search, we employ 8 groups, yielding 8 candidate summaries. The maximum length of the generated candidate summaries is specified at 80 words. Using an identical length guarantees uniformity between summaries in different languages. This uniformity makes the evaluation process easier and allows for fair comparisons. Moreover, languages differ in terms of word length and sentence structure. The summary word restriction guarantees all summaries are reasonably equal in terms of their content and depth, independent of language. For validation and test sets, we also generate the candidate summaries. To encode reference summaries, candidate summaries, and documents, we initiate the evaluation function with the XLM-RoBERTa checkpoint [^8] FacebookAI/xlm-roberta-base as it learns the inner representation of 100 languages. Then, we apply the contrastive learning objective described in Section 3. For training, Adam (Bengio and LeCun, 2015) optimizer is used with learning rate scheduling. The batch size is set to 2 for candidate generation and 4 for model training after trying 32, 16, 8 and 4 because of our GPU limitations. This indicates that we always choose the maximum batch size allowed by our GPU memory. We evaluate the performance of ConVerSum on the validation set at every 1000 steps using the LaSE. We set 15 epochs for training.
@@
4.7. Experimental Results on High-Resource Large Scale Dataset
This section conveys the experimental findings from the analysis of the ConVerSums along with baselines for the high-resource large-scale dataset. The models’ performance is evaluated using standard metrics like LaSE and BERTScore, which serve as a complete assessment of the quality of the generated summaries. There is no significant difference between the training time for the baseline and ConVerSum. The following sections describe the experimental results, providing insights into the proposed methodologies’ effectiveness and reliability for high-resource large-scale dataset.
4.7.1. Model Performance on CNN/DailyMail Test Set
The performance of ConVerSum based on both mT5 and Flan-T5 on the test set of the CNN/DailyMail dataset is shown in Table 3 and Table 4, respectively.
Table 3. Performance of baseline and ConVerSum based on mT5 on the Test set. When we generate candidate summaries using mT5, we choose mT5 as the baseline model to compare our model with it. LaSE, designed for specifically CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
Evaluation
Score
Baseline (mT5)
ConVerSum
on mT5
LaSE 0.3886 0.4408
BERTScore 0.8315 0.8363
Table 4. Performance of baseline and ConVerSum based on Flan-T5 on the Test set. When we generate candidate summaries using Flan-T5, we choose Flan-T5 as the baseline model to compare our model with it. LaSE, designed for specifically CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
Evaluation
Score
Baseline (Flan T5)
ConVerSum
on Flan T5
LaSE 0.3585 0.4012
BERTScore 0.8411 0.8381
In comparison to the mT5 and Flan-T5 baselines, ConVerSum performs better, as demonstrated by the findings shown in Table 3 and Table 4. The ConVerSum’s BERTScore on Flan-T5 is almost identical to the baseline’s, however, the overall performance metrics show a distinct benefit. Furthermore, ConVerSum regularly performs better than the Flan-T5 when compared to the mT5.
4.7.2. Performance Comparison of ConVerSum with Baselines on CrossSum Dataset
@@
Table 5. Performance comparison of (English source and other target) baseline and ConVerSum based on mT5 on CrossSum. When we generate candidate summaries using mT5, we choose mT5 as the baseline model to compare our model with it. LaSE, designed for specifically CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
@@
Key Findings: ConVerSum outperforms the baseline for LaSE, ensuring CLS robustness. Besides, ConVerSum performs better/almost similar to the baseline for BERTScore emphasizing the meaning similarity.
Table 6. Performance comparison of (English source and Others target) baseline and ConVerSum based on Flan-T5 on CrossSum. When we generate candidate summaries using Flan-T5, we choose Flan-T5 as the baseline model to compare our model with it. LaSE, designed for specifically CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
@@
Key Findings: ConVerSum performs better/almost similar to the baseline for LaSE. Besides, ConVerSum performs almost similarly to the baseline for BERTScore. However, it has significantly lower performance than ConVerSum on mT5(mentioned in Table 5) as the generated candidates fail to show diverse quality and languages for Flan-T5.
Table 7. Performance comparison of (Others source and English target) baseline and ConVerSum based on mT5 on CrossSum. When we generate candidate summaries using mT5, we choose mT5 as the baseline model to compare our model with it. LaSE, designed for specifically CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
@@
Key Findings: ConVerSum shows significantly improved performance for Chinese_simplified-English. However, it shows lower than baseline performance for Tigrinya-English. Moreover, our model shows almost similar to baseline performance for Bengali-English as the generated candidates fail to show diverse quality summaries in Tigrinya and Bengali. Besides, ConVerSum outperforms the baseline for the rest of the pairs.
Table 8. Performance comparison of (Others source and English target) baseline and ConVerSum based on Flan-T5 on CrossSum. When we generate candidate summaries using Flan-T5, we choose Flan-T5 as the baseline model to compare our model with it. LaSE, designed specifically for CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
@@
Key Findings: ConVerSum outperforms baseline for LaSE except Bengali-English However, it shows significantly lower performance than ConVerSum on mT5 (mentioned in Table 7) as the generated candidates fail to show diverse quality and languages for Flan-T5.
Table 9. Performance comparison of (Others source and Others target) baseline and ConVerSum based on mT5 on CrossSum. When we generate candidate summaries using mT5, we choose mT5 as the baseline model to compare our model with it. LaSE, designed specifically for CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
@@
Key Findings: ConVerSum shows significantly improved performance for Japanese-Burmese, and Thai-Tamil than the baseline. However, it shows lower than baseline performance for Hindi-Uzbek, and Marathi-Russian as the generated candidates fail to show diverse quality summaries in Hindi-Uzbek, and Marathi-Russian. Otherwise, ConVerSum outperforms the baseline for the rest of the pairs.
Table 10. Performance comparison of (Others source and Others target) baseline and ConVerSum based on Flan-T5 on CrossSum. When we generate candidate summaries using Flan-T5, we choose Flan-T5 as the baseline model to compare our model with it. LaSE, designed specifically for CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
@@
Key Findings: ConVerSum outperforms the baseline for LaSE except for Hindi-Uzbek, and Marathi-Russian. However, it shows significantly lower performance than ConVerSum on mT5 (mentioned in Table 9) as the generated candidates fail to show diverse quality and languages for Flan-T5.
4.8. Second Consideration: Training ConVerSum on Low-Resource Language Small Datasets
This section emphasizes how we develop ConVerSum for CLS on XLSUM (Hasan et al., 2021) Dataset as we mentioned in Section 4.3 to observe how ConVerSum performs on the low-resource small-scale datasets. We choose Bengali, Thai, Burmese and Tigrinya from the XLSUM dataset. As the results using mT5 outperform Flan-T5 in Section 4.7.2, we choose only using mT5 in the case discussed in this section. Therefore, all training setups are kept the same setup mentioned in Section 4.6.
4.9. Experimental Results on Low-Resource Small Scale Dataset
4.9.1. Model Performance on XLSUM Test Set
The performance of ConVerSum based on mT5 for training on four separate low-resource languages - Bengali, Burmese, Thai and Tigrinya on the test set of the XLSUM dataset is shown in Table 11.
Table 11. Performance of baseline and ConVerSum based on mT5 on the Test set of XLSUM for four separate languages-Bengali, Burmese, Thai, Tigrinya. As we generate candidate summaries using mT5, we choose mT5 as the baseline model to compare ConVerSum with it. LaSE, designed for specifically CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
Training Datasets
Evaluation
Metrics
Baseline(mT5) ConVerSum Findings
LaSE 0.5872 0.6228
XLSUM-Bengali BERTScore 0.8658 0.8690
LaSE 0.5579 0.5805
XLSUM-Burmese BERTScore 0.8759 0.8798
LaSE 0.4384 0.5055
XLSUM-Thai BERTScore 0.8475 0.8472
LaSE 0.5265 0.5436
XLSUM-Tigrinya BERTScore 0.8747 0.8768 ConVerSum outperforms baselines for all cases
4.9.2. Performance Comparison of ConVerSum Trained on XLSUM-Bengali with Baseline on CrossSum Dataset
Table 12. Performance Comparison of ConVerSum Trained on XLSUM-Bengali with Baseline on CrossSum Dataset
Baseline (mT5) Proposed Model
Source-Target LaSE BERTScore LaSE BERTScore
Bengali-English 0.5168 0.8263 0.5148 0.8257
Bengali-Arabic 0.5650 0.8229 0.5535 0.8205
Bengali-Marathi 0.6197 0.8493 0.6357 0.8545
Ukrainian-Bengali 0.6110 0.8493 0.6114 0.8470
Hausa-Bengali 0.5831 0.8493 0.6846 0.8629
Key Findings: ConVerSum shows significant improvement in performance for Hausa-Bengali though lower than baseline performance for Bengali-Arabic. However, it shows almost similar to the baseline performance for Bengali-English. The generated candidates fail to show diverse quality summaries for Arabic. ConVerSum outperforms baseline for the rest of the pairs.
Table 13. Performance Comparison of ConVerSum Trained on XLSUM-Thai with Baseline on CrossSum Dataset
Source-Target Baseline(mT5) Proposed Model
LaSE BERTScore LaSE BERTScore
Thai-English 0.5078 0.8298 0.5223 0.8315
Thai-Bengali 0.5937 0.8445 0.6447 0.8487
Thai-Marathi 0.5998 0.8411 0.6677 0.8465
Uzbek-Thai 0.6672 0.9305 0.6672 0.9390
Amharic-Thai 0.3783 0.8340 0.4310 0.8268
Key Findings: ConVerSum outperforms baseline for all of the pairs and shows significant improvement for Thai-Bengali, Thai-Marathi, Amharic-Thai.
Table 14. Performance Comparison of ConVerSum Trained on XLSUM-Burmese with Baseline on CrossSum Dataset
Source-Target Baseline(mT5) Proposed Model
LaSE BERTScore LaSE BERTScore
Burmese-English 0.5394 0.8287 0.5422 0.8283
Burmese-Bengali 0.6146 0.8400 0.5973 0.8448
Burmese-Marathi 0.4247 0.8371 0.6821 0.8648
Uzbek-Burmese 0.4918 0.8335 0.4794 0.8431
Amharic-Burmese 0.6345 0.8689 0.6054 0.8599
Key Findings: ConVerSum shows significant improvement than baseline performance for Burmese-Marathi, both are low-resource languages and also outperform the Burmese-English. Moreover, our model shows better semantically similar performance than the baseline for Burmese-Bengali and Uzbek-Burmese. However, our model performs lower than the baseline only for Amharic-Burmese.
Table 15. Performance Comparison of ConVerSum Trained on XLSUM-Tigrinya with Baselines on CrossSum Dataset
Baseline(mT5) Proposed Model
Source-Target LaSE BERTScore LaSE BERTScore
Tigrinya-English 0.4478 0.7935 0.3758 0.7937
Tigrinya-Arabic 0.4137 0.7837 0.5241 0.7966
Tigrinya-Marathi 0.4185 0.7762 0.6423 0.7916
Chinese_simplified-Tigrinya 0.4390 0.8288 0.5497 0.8288
Somali-Tigrinya 0.2241 0.7829 0.5537 0.7801
Key Findings: ConVerSum shows significant improvement over baseline performance for all pairs except Tigrinya-English.
4.10. Performance Comparison of ConVerSum and LLMs for Low Resource Languages by Prompting
@@
Evaluating the performance of the state-of-art LLMs on various datasets and numerous NLP tasks is a trending research (Laskar et al., 2023; Kabir et al., 2023). Inspired by this trend, in this section, we compare the performance of ConVerSum to cutting-edge LLMs - ChatGPT-3.5, ChatGPT-4o, and Gemini 1.5 Pro to measure ConVerSum’s efficiency in generating high-quality summaries in a variety of low-resource languages. To assess LLMs’ proficiency, we prompt them to rate their level of confidence between 1 to 10 in generating accurate cross-lingual summaries in 45 different languages. Responses are repeatedly generated to ensure clarity. All LLMs rate their capability as low (1-4) for CLS on low-resource languages. We choose three language pairs mentioned in Table 16 based on their ratings for this experiment.
@@
”How confident are you to generate high-quality cross-lingual summary concisely and informatively for low-resource languages? Find the list of languages - Amharic, Arabic, Azerbaijani, Bengali, Burmese, Chinese_simplified, Chinese_traditional, English, French, Gujarati, Hausa, Hindi, Igbo, Indonesian, Japanese, Kirundi, Korean, Kyrgyz, Marathi, Nepali, Oromo, Pashto, Persian, Pidgin, Portuguese, Punjabi, Russian, Scottish_gaelic, Serbian_cyrillic, Serbian_latin, Sinhala, Somali, Spanish, Swahili, Tamil, Telugu, Thai, Tigrinya, Turkish, Ukrainian, Urdu, Uzbek, Vietnamese, Welsh, Yoruba. Rate your confidence level for cross-lingual summarization on a scale of 1 to 10 for the given languages.”
@@
Then, we use manual prompting for LLMs which is free of cost. For our selected language pairs, we provide the following prompt where the source document is provided from CrossSum (Bhattacharjee et al., 2023) test set and extract the responses manually from LLMs to measure the LaSE and BERTScore and compare those with ConVerSum which are mentioned in Table 16.
”Summarize the given text in target language, preferably in 80 words, concisely and informative. .. (source document is given here)”
Table 16. Some Performance Measures on LLM vs ConVerSum for Low-resource Languages
@@
4.10.1. Limitations of Prompt-based Approach
Although this manual prompting and extracting approach is budget-friendly, it is very time-consuming for broad-level experiments. Moreover, this approach limits handling long documents. Besides, it has confined us to conducting thorough experiments. For this reason, we experiment with using API to conduct a thorough experiment discussed in Section 4.11. We only choose GPT 4o latest version model as it is a more powerful LLM than GPT 3.5 or Gemini.
4.11. Performance Comparison of GPT-4o using API and ConVerSum
This section offers a comparison of ConVerSum with GPT-4o for a range of language pairings in a zero-shot setup. Zero-shot learning refers to the model’s capacity to provide summaries for language pairings that were not explicitly included in its training data. For this purpose, we use the GPT-4o model version gpt-4o-2024-05-13101010https://platform.openai.com/docs/models/gpt-4o, which is the most recent version and represents the model’s status as of October 2023. This version can handle up to 128,000 tokens and a maximum context window of 4,096 tokens. Furthermore, we want to explore how GPT 4o performs on a one-shot setup which is a form of machine learning in which a model learns to recognize objects or execute tasks based on only one training sample per class.
Table 17. Performance comparison of ConVerSum and GPT 4o in zero-shot setup using API
GPT 4o (current version) ConVerSum
Source-Target No of Samples LaSE BERTScore LaSE BERTScore
Burmese-Bengali 14 0.7272 0.7995 0.6030 0.8452
Amharic-Igbo 3 0.7142 0.7370 0.6377 0.7836
Tigrinya-Arabic 2 0.5903 0.7538 0.5241 0.7936
Japanese-Burmese 14 0.5620 0.7853 0.6500 0.8613
Thai-Tamil 40 0.6784 0.7943 0.6445 0.8673
Hausa-Bengali 23 0.6532 0.7759 0.6846 0.8629
Chinese_simplified-Tigrinya 1 0.3178 0.7151 0.5497 0.8288
Burmese-Marathi 9 0.6606 0.7912 0.6821 0.8648
Thai-Marathi 20 0.6743 0.7937 0.6677 0.8465
Chinese_simplified-English 486 0.8156 0.8162 0.5361 0.8437
Chinese_traditional-Ukrainian 143 0.6520 0.8022 0.5624 0.8305
Uzbek-Thai 3 0.5857 0.7173 0.6672 0.9390
4.11.1. Performance Analysis on Zero-Shot Setup
Table 17 illustrates the performance comparison between GPT-4o and ConVerSum across various language pairs. We conduct experiments on 12 separate language pairs. We consider several combinations for choosing language pairs as per GPT 4o’s confidence level such as low-low, high-low, high-high, mid-mid, and so on.
Among 12 language pairs, ConVerSum outperforms 5 language pairs GPT 4o in which 3 pairs are not included in the list of the high confidence level of CLS for GPT-4o. and 2 pairs have at least one between each source and target pair from a high confidence level for GPT-4o. ConVerSum shows almost similar performance for 1 pair where both source and target are from mid-confident languages by GPT 4o.
ConVerSum shows a clear benefit in dealing with low-resource languages, as observed in pairs like Burmese-Marathi and Uzbek-Thai, Hausa-Bengali which pairs are not included in list of the high confidence level of CLS for GPT-4o.
There are also two language pairs in which ConVerSum outperforms GPT 4o - Japanese-Burmese and Chinese_simplified-Tigrinya. Among these languages, we find at least one between each source and target pairs such as Japanese and Chinese_simplified from the high confidence languages for GPT 4o.
This is seen by language pairs like Chinese_simplified-English, where it received the highest LaSE score of 0.8156, indicating stronger generalization for high-resource languages. ConVerSum performs significantly lower than GPT 4o for source high-confident and target high-confident language pair combinations.
The proposed model’s constant outperformance in BERTScore, despite variations in LaSE scores, indicates its power in capturing semantic and contextual complexities, especially in language pairings with limited resources where GPT-4o’s performance is somewhat weaker. From this discussion, we conclude that in situations when data availability is constrained, how well the model can perform CLS tasks.
4.11.2. Performance Analysis on One-Shot Setup
We want to explore the performance of the GPT 4o model in a one-shot setup. For this purpose, we choose 3 low-resource language pairs. We provide only one example from the train set of the CrossSum dataset from the respective language-pair datasets. Table 18 illustrates the performance comparison among GPT 4o’s zero-shot, one-shot and ConVerSum.
Table 18. Performance comparison of ConVerSum and GPT 4o in zero-shot and one-shot setup using API
GPT 4o (current model)
Zero Shot One shot ConVerSum
Source-Target No of Samples LaSE BERTScore LaSE BERTScore LaSE BERTScore
Burmese-Bengali 14 0.7272 0.7995 0.7204 0.8001 0.6030 0.8452
Amharic-Igbo 3 0.7142 0.7370 0.6871 0.7352 0.6377 0.7836
Tigrinya-Arabic 2 0.5903 0.7538 0.5222 0.7551 0.5241 0.7936
From Table 18, we observe that the one-shot performance of GPT 4o for LaSE has been decreased than zero-shot for GPT 4o. As CLS can be challenging to LLMs when there’s not enough contextual understanding or fine-tuning for the specific language pairs, they struggle to perform well when it involves nuances specific to certain languages like idioms or syntax. For 1 language pair (Tigrinya-Arabic), ConVerSum outperforms GPT 4o in one-shot setup. ConVerSum outperforms BERTScore for all pairs we experimented with highlighting capturing semantic and contextual complexities.
4.12. Discussion on LLM for CLS